Give scientific reasons:(1) Rate of reaction is directly proportional to concentration of reactant.(2)Rate of reaction becomes double for every 10K rise in temperature.
1
Expert's answer
2017-03-07T05:05:49-0500
Answer:
a) According to collision theory, for reaction to take place, reacting molecules must collide. Chemical bounds are formed as a result of those collisions. Therefore, as the concentration of the reactant increases, the number of possible chemical bounds also rises. As a result, reaction rate also increases.
b) As the temperature rises, energy of each molecule increases. Therefore, molecule’s collisions more possibly lead to the formation of chemical bounds. According to the Arrhenius theory, number of molecules having this minimum energy is found to be proportional to e-Ea/RT.
k = k0 · exp(-Ea/RT)
k1/k2 = exp(-Ea/R · (1/T2 – 1/T1))
Putting, T1=298K, T2=308K and Ea = 53 kJ/mol
k2/k1=2